Shift report 20011230 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Dec 30 2001 - 23:47:32 EST

  • Next message: Apache: "Shift report 20011231 00:00-08:00"

    ------------------------------------------------------------
    Date: Sun Dec 30 23:47:32 2001, Shift: 16:00-24:00
    Supervisor: R. Debbe, with: 
    Log book#11 pages 123 - 125
    
    Shift Summary:
    ------------------------------------------------------------
    1. One good store that started at 17:41 (the one from the previous shift was dumped at 16:30). At 19:50 a sextuple quench dumped the beam, since then, no beam. At 23:30 both rings were filled and ready for ramping but the beam was dumped just after transition by a Quench Link Interlock
    
    2. I collected 90K trigger 6 in run 6215, we now have some 140K trig 6 at 1/4 A at 8 degrees
    
    3. No detector problems
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Dec 30 2001 - 23:47:59 EST